RECOVERY OF LEGACY PAPER SPECTRA AND NEW RESULTS ON THE ROTATIONAL SPECTRUM OF H_{2}O⋯HF

\underline{ZBIGNIEW KISIEL}, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; MIKHAIL YU. TRETYAKOV, OLEG L. POLYANSKY, Institute of Applied Physics of RAS, 46 Uljanova Str., 603950 Nizhny Novgorod, Russia.

The situation when a spectrum exists only in the form of a paper record and contains valuable unprocessed information is not that rare. In the case of broadband spectra with a rich information content it is very desirable to convert such spectra into a digital form that will be amenable for use with contemporary packages for graphical assignment.^{a}

The RAD spectrum of the H_{2}O⋯HF hydrogen bonded dimer recorded in the 1980’s at Nizhny Novgorod is an example of such a valuable legacy spectrum. Only a small proportion of the lines in this spectrum have been assigned and measured.^{b} A special software package has been written to convert the original spectrum from many strip chart records into a unified, calibrated digital spectrum, which covers the frequency region 182-352 GHz and 11 succesive J + 1 ← J transitions. The package is described in detail, is expected to be of universal applicability, and has been made available on the PROSPE website.^{c} Initial results concerning the analysis of this spectrum are reported, and it is aimed to considerably improve the understanding of the internal dynamics in H_{2}O⋯HF over that reached earlier on the basis of the cm-wave rotational spectrum.^{d}

aZ.Kisiel et al., J. Mol. Spectrosc., 233, 231-243 (2005).

bS.P.Belov et al., J. Mol. Spectrosc., 241, 124-135 (2007).

cZ.Kisiel, Programs for ROtational SPEctroscopy, available at http://info.ifpan.edu.pl/˜ kisiel/prospe.htm.

dZ.Kisiel, A.C.Legon, D.J.Millen, Proc. Roy. Soc. Lond. A, 381, 419-442 (1982).